OpenStudy (anonymous):

the shorter leg of a 30 degree, -60 degree, - 90 degree triangle is 6 what is the length of the other leg ,Special Right Triangle

OpenStudy (angela210793):

The side in front of 30 degree angle is always Hypotenuse/2(in a right triangle) so hypotenuse is 2*6=12 and the missing side u can find it by using Pythagorean theorem x^2=12^2-6^2 x^2=144-36 X=sqr108=6sqr3